# Tiny Time Mixers (TTMs): Fast Pre-trained Models for Enhanced Zero/Few-Shot Forecasting of Multivariate Time Series

### v7

@@ Abstract
Large pre-trained models excel in zero/few-shot learning for language and vision tasks but face challenges in multivariate time series (TS) forecasting due to diverse data characteristics. Consequently, recent research efforts have focused on developing pre-trained TS forecasting models. These models, whether built from scratch or adapted from large language models (LLMs), excel in zero/few-shot forecasting tasks. However, they are limited by slow performance, high computational demands, and neglect of cross-channel and exogenous correlations. To address this, we introduce Tiny Time Mixers (TTM), a compact model (starting from 1M parameters) with effective transfer learning capabilities, trained exclusively on public TS datasets. TTM, based on the light-weight TSMixer architecture, incorporates innovations like adaptive patching, diverse resolution sampling, and resolution prefix tuning to handle pre-training on varied dataset resolutions with minimal model capacity. Additionally, it employs multi-level modeling to capture channel correlations and infuse exogenous signals during fine-tuning. TTM outperforms existing popular benchmarks in zero/few-shot forecasting by (4-40%), while reducing computational requirements significantly. Moreover, TTMs are lightweight and can be executed even on CPU-only machines, enhancing usability and fostering wider adoption in resource-constrained environments. Model weights for our initial variant TTM are available here. Model weights for more sophisticated variants (TTM, TTM, and TTM) will be shared soon. The source code for TTM can be accessed here.
@@ 4.4 TTM’s Zero-shot Performance and Inference Cost
Recently, popular pre-trained models like TimesFM, Moirai, Chornos, Lag-llama, and LLMTime have gained traction for their zero-shot (ZS) forecasting capabilities. Among these, Chornos, Lag-llama, and LLMTime suffer from lengthy ZS inference time, posing practical challenges for testing across all sliding windows of the test set. To address this, LLMTime suggests using the last test window for benchmarking, a practice we also adopt for comparing with this set of SOTA models. On the other hand, TimesFM and Moirai exhibit comparatively faster ZS inference speeds, enabling testing across all sliding windows of the test set. Table 1 presents a comparison of TTM performance with Moirai and TimesFM. Despite having significantly fewer parameters, the variants of TTM consistently outperform most benchmark variants. Notably, TTM, which is 3-62X smaller than all Moirai variants and 40X smaller than TimesFM, outperforms the Moirai variants by 4-10% and TimesFM by 19%. Even TTM, with just 1M parameters, outperforms most benchmarks by a considerable margin, highlighting the effectiveness of TTM. Moreover, as depicted in Appendix E.4, TTM zero-shot results consistently outperform the full-shot results of popular architectures in short context length settings. Likewise, Table 2 presents a comparison of TTM performance with Chronos and Lag-llama on the last test-window set. As indicated, TTM which is 8-709X smaller than Chronos, outperforms it by 17-32%. Likewise TTM, which is 2-3X smaller than Lag-llama, outperforms it by 40%. In addition, TTM also outperforms the massive LLMTime and UniTime by over 25% as reported in Appendix E.3. Table 3 presents the inference time per batch and maximum GPU memory requirement of different TS pre-trained models. Notably, TTM exhibits the lowest inference time and memory usage among them.
@@ 4.5 TTM’s Few-shot and Full-shot Head Probing Performance
In operational deployments, users typically leverage a small set of target data for fine-tuning to enhance the model performance. In this regard, TTM provides a highly efficient quick fine-tuning process, enabling users to enhance forecasting accuracy swiftly by training only the model head. GPT4TS and Time-LLM are two state-of-the-art pre-trained models that present results for few-shot training. As demonstrated in Table 4, TTM surpasses GPT4TS by 15% and Time-LLM by 10% in the few-shot 5% setting, where only 5% of the train data is used for fine-tuning. In addition, we also report the Few-shot 5% results of several popular SOTA architectures in Table 4, where TTM demonstrates superior performance. This underscores the significance of TTM’s pre-trained weights, which substantially contribute to its effectiveness in data-constrained scenarios. Likewise, TTM also excels in few-shot cross-transfer learning tasks outperforming popular SOTAs (including SimMTM [5]) as shown in the Appendix E.6.
Alternatively, if the train split of the complete target dataset is available, head probing using the entire dataset becomes feasible. This involves fine-tuning the model head using all available data while keeping the backbone weights unchanged. Recently, the Moment [10] model has achieved the SOTA results in head probing as compared to GPT4TS and Time-LLM. However, as indicated in Table 5, TTM further outperforms the results reported by Moment by 3-4%. In addition, TTM head probing results are very competitive as compared to the full end-to-end training of popular architectures as depicted in Appendix E.7. Hence, TTM, with its significantly reduced model size and the absence of compute-intensive components like self-attention, enables quick fine-tuning of models compared to the cumbersome process required by the massive Transformer models. Note that Moment is excluded from the comparison of zero/few-shot forecasting results because it does not report them.
@@
| Models | BS | CC | APP | SER | f-imp% | |
| Zero-shot | TTM | 0.992 | 0.263 | 0.183 | 0.238 | 44% |
| Moirai | 1.034 | 0.292 | 0.245 | 0.292 | 50% | |
| Moirai | 0.885 | 0.286 | 0.203 | 0.264 | 46% | |
| Moirai | 0.73 | 0.286 | 0.214 | 0.261 | 42% | |
| TimesFM | 0.668 | 0.321 | 0.274 | 0.297 | 45% | |
| Fine-Tune | TTM | 0.635 | 0.261 | 0.073 | 0.143 | 18% |
| PatchTST | 0.735 | 0.267 | 0.060 | 0.119 | 15% | |
| TSMixer-CC | 0.651 | 0.284 | 0.053 | 0.136 | 15% | |
| TSMixer-CM | 0.716 | 0.303 | 0.069 | 0.118 | 20% | |
| TSMixer | 0.664 | 0.267 | 0.066 | 0.134 | 17% | |
| GPT4TS | 0.645 | 0.254 | 0.075 | 0.135 | 18% | |
@@ 4.6 TTM’s Effectiveness in Cross-channel and Exogenous Modeling
Since the datasets (D1) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets (D2, as explained in Section 4.1) to quantify its benefits. Since these datasets are already very small, we used their full data for fine-tuning. Table 6 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare TTM-CM with plain TTM finetuning and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with cross-channel reconciliation head (TSMixer-CC) [6] as they are the latest SOTAs in channel-correlation modelling. In addition, we also report the zero-shot numbers of the best pre-trained models for reference. From Table 6, we can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
@@ 4.7 Ablation Studies
Effect of Resolution Prefix Tuning (RPT) and Adaptive Patching(AP): RPT enhances forecast performance, especially with large and diverse pretraining (PT) data. Adding a learnable resolution prefix token allows models to easily decouple weights across different resolutions, leading to a 3% improvement in 1B PT data setup (Table 7). RPT is also beneficial for very short context length scenarios, improving performance by 8% (Appendix E.9). On the other hand, AP generally improves the forecasting performance across all set-ups, but the impact is consistently high in less PT data settings (3% boost). Further details are in Appendix E.8.
@@ 4.8 TTM Model Insights & Explainability
Figure 5 illustrates the TTM embeddings from various datasets (weather, traffic, and electricity) using PCA projection, each represented by a different color. From each dataset, three distant, non-overlapping, fixed-length time segments (S-1, S-2, S-3) are selected, each depicted with a unique marker shape. The visualization uses the first and second principal components of the TTM embeddings. The inset image focuses on the weather dataset alone, revealing a deeper structure learned by the TTM architecture. The cyclic orbits in the embeddings reflect the seasonal patterns in the data. Both hourly datasets (traffic and electricity) form concentric orbits due to similar seasonal patterns, while the weather data, with its distinct seasonal pattern, shows cyclic orbits in a different sub-dimension. In addition, the cross-channel attention from the fine-tuned model’s channel mixing layers reveals feature importance across channels. As shown in Figure 5, the model focuses on channels like weathersit, season, holiday, and temperature to predict bike-rental counts. These attention model weights correlate with the general data characteristics where bike rental demands are heavily influenced by weather and holidays, providing explanation for the fine-tuned model predictions. More details are in Appendix F.
@@ C.1 List of Pre-training Datasets
Pre-training employs a subset of 1B samples from Monash [9] and Libcity [32, 35] data collection, where Monash results in 250M samples and LibCity accounts for the rest. In this estimate, one sample denotes a pair of training windows: and . We employ a subset of the datasets available in the Monash forecasting data repository [9] available at https://forecastingdata.org/. Since our primary focus in this study is long term forecasting with forecast length ranging from 96 to 720, it is not possible to use yearly, monthly, quarterly, or weekly datasets due to their short lengths. Hence, we skip a few datasets of short lengths. The Monash datasets used are available under a Creative Commons Attribution 4.0 International license. For LibCity, we employ all datasets released by the Moirai authors [35], available at https://huggingface.co/datasets/Salesforce/lotsa_data/tree/main (except the Rotterdam dataset which was not available during our experimentation). The LibCity datasets at the above link were released under an Apache 2.0 license. The final list of all pre-training datasets is shown in Table 8.
@@ D.4 Baseline Implementation Details
Appendix E Full Results Tables
@@
E.1 Full table for all TTM variants
E.2 Full table for zero-shot experiment
@@
E.3 Other zero-shot comparisons
@@
E.4 TTM Zero-shot vs. SOTA Full-shot (short context setting)
@@
E.5 Full table for 5% few-shot experiment
@@
E.6 TTM vs. Cross-transfer models
@@
E.7 Full table for full-shot head-probing experiment
@@
E.8 Full table: Impact of Adaptive Patching (AP)
@@
E.9 Full table: Impact of Resolution Prefix Tuning (RPT)
@@
Appendix F Model Insights and Explanation
F.1 Dataset preparation for TTM embedding analysis
@@ F.2 Channel Attention Map
Appendix G Limitations and Future Work

### v8

@@ Abstract
Large pre-trained models excel in zero/few-shot learning for language and vision tasks but face challenges in multivariate time series (TS) forecasting due to diverse data characteristics. Consequently, recent research efforts have focused on developing pre-trained TS forecasting models. These models, whether built from scratch or adapted from large language models (LLMs), excel in zero/few-shot forecasting tasks. However, they are limited by slow performance, high computational demands, and neglect of cross-channel and exogenous correlations. To address this, we introduce Tiny Time Mixers (TTM), a compact model (starting from 1M parameters) with effective transfer learning capabilities, trained exclusively on public TS datasets. TTM, based on the light-weight TSMixer architecture, incorporates innovations like adaptive patching, diverse resolution sampling, and resolution prefix tuning to handle pre-training on varied dataset resolutions with minimal model capacity. Additionally, it employs multi-level modeling to capture channel correlations and infuse exogenous signals during fine-tuning. TTM outperforms existing popular benchmarks in zero/few-shot forecasting by (4-40%), while reducing computational requirements significantly. Moreover, TTMs are lightweight and can be executed even on CPU-only machines, enhancing usability and fostering wider adoption in resource-constrained environments. The model weights for reproducibility and research use are available here, while enterprise-use weights under the Apache license can be accessed as follows: the initial TTMvariant here, and the latest variants (TTM, TTM, TTM) weights are available here (preferred use). The source code for the TTM model along with the usage scripts are available here.
[Accepted at the 38th Conference on Neural Information Processing Systems (NeurIPS 2024). This work is submitted to arXiv on Jan 08, 2024.]
@@ 4.4 TTM’s Zero-shot Performance and Inference Cost
Recently, popular pre-trained models like TimesFM, Moirai, Chornos, Lag-llama, and LLMTime have gained traction for their zero-shot (ZS) forecasting capabilities. Among these, Chornos, Lag-llama, and LLMTime suffer from lengthy ZS inference time, posing practical challenges for testing across all sliding windows of the test set. To address this, LLMTime suggests using the last test window for benchmarking, a practice we also adopt for comparing with this set of SOTA models. On the other hand, TimesFM and Moirai exhibit comparatively faster ZS inference speeds, enabling testing across all sliding windows of the test set. Table 1 presents a comparison of TTM performance with Moirai and TimesFM. Despite having significantly fewer parameters, the variants of TTM consistently outperform most benchmark variants. Notably, TTM, which is 3-62X smaller than all Moirai variants and 40X smaller than TimesFM, outperforms the Moirai variants by 4-10% and TimesFM by 19%. Even TTM, with just 1M parameters, outperforms most benchmarks by a considerable margin, highlighting the effectiveness of TTM. Moreover, as depicted in Appendix F.4, TTM zero-shot results consistently outperform the full-shot results of popular architectures in short context length settings. Likewise, Table 2 presents a comparison of TTM performance with Chronos and Lag-llama on the last test-window set. As indicated, TTM which is 8-709X smaller than Chronos, outperforms it by 17-32%. Likewise TTM, which is 2-3X smaller than Lag-llama, outperforms it by 40%. In addition, TTM also outperforms the massive LLMTime and UniTime by over 25% as reported in Appendix F.3. Table 3 presents the inference time per batch and maximum GPU memory requirement of different TS pre-trained models. Notably, TTM exhibits the lowest inference time and memory usage among them.
@@ 4.5 TTM’s Few-shot and Full-shot Head Probing Performance
In operational deployments, users typically leverage a small set of target data for fine-tuning to enhance the model performance. In this regard, TTM provides a highly efficient quick fine-tuning process, enabling users to enhance forecasting accuracy swiftly by training only the model head. GPT4TS and Time-LLM are two state-of-the-art pre-trained models that present results for few-shot training. As demonstrated in Table 4, TTM surpasses GPT4TS by 15% and Time-LLM by 10% in the few-shot 5% setting, where only 5% of the train data is used for fine-tuning. In addition, we also report the Few-shot 5% results of several popular SOTA architectures in Table 4, where TTM demonstrates superior performance. This underscores the significance of TTM’s pre-trained weights, which substantially contribute to its effectiveness in data-constrained scenarios. Likewise, TTM also excels in few-shot cross-transfer learning tasks outperforming popular SOTAs (including SimMTM [5]) as shown in the Appendix F.6.
Alternatively, if the train split of the complete target dataset is available, head probing using the entire dataset becomes feasible. This involves fine-tuning the model head using all available data while keeping the backbone weights unchanged. Recently, the Moment [10] model has achieved the SOTA results in head probing as compared to GPT4TS and Time-LLM. However, as indicated in Table 5, TTM further outperforms the results reported by Moment by 3-4%. In addition, TTM head probing results are very competitive as compared to the full end-to-end training of popular architectures as depicted in Appendix F.7. Hence, TTM, with its significantly reduced model size and the absence of compute-intensive components like self-attention, enables quick fine-tuning of models compared to the cumbersome process required by the massive Transformer models. Note that Moment is excluded from the comparison of zero/few-shot forecasting results because it does not report them.
@@
| Models | BS | CC | APP | SER | f-imp% |
| TTM | 0.635 | 0.261 | 0.073 | 0.143 | 18% |
| PatchTST | 0.735 | 0.267 | 0.060 | 0.119 | 15% |
| TSMixer-CC | 0.651 | 0.284 | 0.053 | 0.136 | 15% |
| TSMixer-CM | 0.716 | 0.303 | 0.069 | 0.118 | 20% |
| TSMixer | 0.664 | 0.267 | 0.066 | 0.134 | 17% |
| GPT4TS | 0.645 | 0.254 | 0.075 | 0.135 | 18% |
@@ 4.6 TTM’s Effectiveness in Cross-channel and Exogenous Modeling
Since the datasets (D1) used in previous experiments do not have exogenous variables, we evaluate the effectiveness of TTM on 4 other datasets (D2, as explained in Section 4.1) to quantify its benefits. Since these datasets are already very small, we used their full data for fine-tuning. Table 6 shows the performance of the pre-trained TTM model fine-tuned on the target data with exogenous mixer module and decoder channel-mixing enabled (TTM-CM). We compare TTM-CM with plain TTM finetuning and other primary SOTAs (PatchTST, TSMixer variants, and GPT4TS) trained from scratch. Specifically, we compare with TSMixer with channel-mixing enabled (TSMixer-CM) and TSMixer with cross-channel reconciliation head (TSMixer-CC) [6] as they are the latest SOTAs in channel-correlation modelling. From Table 6, we can see that TTM-CM outperforms all the competitive models with a significant margin (15-44%), thus, demonstrating the power of TTM in capturing inter-channel correlations.
@@ 4.7 Ablation Studies
Effect of Resolution Prefix Tuning (RPT) and Adaptive Patching(AP): RPT enhances forecast performance, especially with large and diverse pretraining (PT) data. Adding a learnable resolution prefix token allows models to easily decouple weights across different resolutions, leading to a 3% improvement in 1B PT data setup (Table 7). RPT is also beneficial for very short context length scenarios, improving performance by 8% (Appendix F.9). On the other hand, AP generally improves the forecasting performance across all set-ups, but the impact is consistently high in less PT data settings (3% boost). Further details are in Appendix F.8.
@@ 4.8 TTM Model Insights & Explainability
Figure 5 illustrates the TTM embeddings from various datasets (weather, traffic, and electricity) using PCA projection, each represented by a different color. From each dataset, three distant, non-overlapping, fixed-length time segments (S-1, S-2, S-3) are selected, each depicted with a unique marker shape. The visualization uses the first and second principal components of the TTM embeddings. The inset image focuses on the weather dataset alone, revealing a deeper structure learned by the TTM architecture. The cyclic orbits in the embeddings reflect the seasonal patterns in the data. Both hourly datasets (traffic and electricity) form concentric orbits due to similar seasonal patterns, while the weather data, with its distinct seasonal pattern, shows cyclic orbits in a different sub-dimension. In addition, the cross-channel attention from the fine-tuned model’s channel mixing layers reveals feature importance across channels. As shown in Figure 5, the model focuses on channels like weathersit, season, holiday, and temperature to predict bike-rental counts. These attention model weights correlate with the general data characteristics where bike rental demands are heavily influenced by weather and holidays, providing explanation for the fine-tuned model predictions. More details are in Appendix G.
4.9 Discussion on TTM Design choices
In this section, we intuitively explain the important design choices of TTM that greatly enhance its forecasting accuracy and transfer learning capabilities despite its extremely small model capacity:
All existing pre-trained models use a very high volume of pretraining data (for example, TimesFM used 300B and Moirai used 27B time-points), hence they naturally require massive model sizes. However, as shown in Figure.4, we observe that “limited” pretraining data with “high resolution diversity” greatly helps in time-series model generalization, as opposed to simply increasing the pretraining data size. This is an important observation and finding that resolution diversity in pretraining data is very crucial for time-series FMs. Based on these findings, we proceed with a well-reduced dataset (1B samples) with high resolution diversity which naturally reduces our model size compared to counterparts needing to pre-train with several hundred billion time-series. We introduce a high diversity in our data via Diverse Resolution Sampling technique (DRS) which our counterparts fail to do.
Secondly, we opted for TSMixer-based models instead of transformer-based models, which further reduced the model size drastically. The TSMixer architecture has successfully established in the past that interleaving simple gated attentions with mixing components across patches, channels, and features greatly enhances forecasting accuracies with very limited model capacity, as the quadratic time-complexity of self-attentions can be entirely avoided. Following TSMixer, several other mixer architectures [33] [24] have been published, reiterating the power of these simple architectures. Thus, avoiding complex transformer architectures further reduced our model size significantly.
In addition, we further increased the modeling power of TSMixer without drastically increasing its size by introducing several innovative components, such as adaptive patching, diverse resolution sampling, and resolution prefix tuning. These enhancements are crucial for effectively handling large pre-training across datasets with varying resolutions, all while keeping the model capacity very minimal.
Finally, framing the pre-training objective as a direct forecasting task demonstrates improved zero-shot performance as compared to the traditional masking-based pre-training approaches. We hypothesize that this method enables the model to effectively learn complex non-linear mappings between the fixed context and forecast windows during pre-training that generalizes well to unseen datasets.
@@ C.1 List of Pre-training Datasets
Pre-training employs a subset of 1B samples from Monash [9] and Libcity [32, 35] data collection, where Monash results in 250M samples and LibCity accounts for the rest. In this estimate, one sample denotes a pair of training windows: and . We employ a subset of the datasets available in the Monash forecasting data repository [9] available at https://forecastingdata.org/. Since our primary focus in this study is long term forecasting with forecast length ranging from 96 to 720, it is not possible to use yearly, monthly, quarterly, or weekly datasets due to their short lengths. Hence, we skip a few datasets of short lengths. The Monash datasets used are available under a Creative Commons Attribution 4.0 International license. For LibCity, we employ all datasets released by the Moirai authors [35], available at https://huggingface.co/datasets/Salesforce/lotsa_data/tree/main (except the Rotterdam dataset which was not available during our experimentation). The LibCity datasets at the above link were released under an Apache 2.0 license. The final list of all pre-training datasets is shown in Table 8. Please note that, the last three datasets in the Libcity section have been excluded from the pre-training process for the model releases intended for enterprise-use.
@@ D.4 Baseline Implementation Details
Appendix E Sample Zero-shot Visualizations
Figure. 6 visualizes the zero-shot forecasts of TTM across different datasets illustrating the power of TTM to capture complex trends and seasonal patterns.
Appendix F Full Results Tables
@@
F.1 Full table for all TTM variants
F.2 Full table for zero-shot experiment
@@
F.3 Other zero-shot comparisons
@@
F.4 TTM Zero-shot vs. SOTA Full-shot (short context setting)
@@
F.5 Full table for 5% few-shot experiment
@@
F.6 TTM vs. Cross-transfer models
@@
F.7 Full table for full-shot head-probing experiment
@@
F.8 Full table: Impact of Adaptive Patching (AP)
@@
F.9 Full table: Impact of Resolution Prefix Tuning (RPT)
@@
Appendix G Model Insights and Explanation
G.1 Dataset preparation for TTM embedding analysis
@@ G.2 Channel Attention Map
Appendix H Limitations and Future Work